### v2

@@ 1 Introduction
Statistical tests for causal inference are ubiquitous, from drug discovery [53] and psychology [61] to social studies and online platforms—being at the core of decision-making. A prevalent example is A/B testing [18], the de facto way to evaluate the potential of a system change before applying it in scenarios such as e-commerce. Typically, the population is split into two random groups, and the treatment is assigned to one of them ($T=1$). The difference between their response variable $Y$ (e.g., time spent on the system) quantifies the average treatment effect (ATE) of the randomized control trial (RCT) [43], i.e., $Y_{T}-Y_{C}=\mathbb{E}[Y|T=1]-\mathbb{E}[Y|T=0]$, where $T$ and $C$ refer to the treatment and control groups, respectively. In this setting, however, we risk causing churn on a massive scale if the proposed change is not effective [2]. It is thus more wise to test on a smaller scale and try to extrapolate our findings. In these cases, we can build a model that predicts the outcome without actually intervening, based on the samples’ confounders, e.g., for a user, this could be the purchase history and demographics. The core hypothesis is that given the common assumption of unconfoundedness [9], samples with similar confounders will exhibit similar outcomes under the same treatment. Developing a model that can predict the effect of a treatment on unseen samples is referred to as uplift modeling (UM) [41], mainly stemming from business context applications. The uplift refers to the prediction of the outcome’s change due to the treatment, i.e., the ATE, for a set of samples [11]. The problem is akin to individual treatment effect (ITE) estimation, i.e., predicting the outcome of the same sample with and without treatment (counterfactual) and computing their difference.
@@ 3.1 Uplift Modeling with Graph Neural Networks (UMGNet)
$\text{{loss}}_{t}=\frac{1}{n}\sum_{i=0}^{n}\text{CrossEntropy}\big{(}T,\text{Sigmoid}(\mathbf{H}^{\top}_{F}\mathbf{W}^{\top}_{F})\big{)}.$ (8)
@@ 3.2 Active Learning for Uplift GNNs (UMGNet-AL)
Aleatoric uncertainty can be measured using the structure of the graph and the feature distance. Acquisition functions based on diverse criteria, including both types of uncertainties, have proven more effective in node-level tasks [62, 16]. We adopt a similar approach and define $D_{u},\forall u\in U$ to be the degree, $Q_{u}$ is the model’s uncertainty, and $b$ is the budget for new training samples in each iteration. According to the literature [62, 16], we define diversity based on feature clustering. Specifically, we cluster the samples with a $k$-means algorithm in a predefined number of clusters and store the assignments in $C\in\mathbb{Z}^{n}$. Then we calculate the distance $M_{u}\in\mathbb{R}^{n}$ between each sample and its cluster centroid and compute the budget for each cluster based on its relative size and $b$, in $C_{b}\in\mathbb{Z}^{|C|}$. We can cast the problem of choosing a batch to add to the training set as ranking based on $U$, $D$, and $M$, with the first constraint being on the batch size. The second constraint, which represents diversity, pertains to how many samples of a given cluster should be included in the train set in each round. Finally, we want to keep the train set balanced in terms of treatment and control samples, hence the third constraint:
@@
The problem can be solved greedily in each iteration of batch selection, while $D$, $C$, and $C_{b}$ are precomputed. In practice, we weigh each of the three criteria in the objective function with a coefficient in $[0,1]$ based on the results of validation.
@@ Appendix 0.A Network as a confounder and as a source for interference
Generally, we can make a distinction between the methods that predict the ATE i) using the network as a confounder [56, 19, 34, 6, 25, 13], ii) assuming interference bias caused by the network [36, 7, 32], iii) separating between confounding and interference bias [35, 8]. That said, all cases are predominantly evaluated based on the predicted ITE’s accuracy, meaning that the interference models do not address how to use the predictions to alleviate the bias from the observed outcome. Our work falls in the first category and follows the same assumptions as the literature regarding interference [56, 34, 6, 13]. Interference in e-commerce bipartite networks is not studied as extensively as in social networks [54, 27]. Its existence remains unproved by experimental studies, mainly because it relies on the type of recommendation algorithm and the nature of the network.

### v4

@@ 1 Introduction
Statistical tests for causal inference are ubiquitous, from drug discovery [54] and psychology [62] to social studies and online platforms—being at the core of decision-making. A prevalent example is A/B testing [19], the de facto way to evaluate the potential of a system change before applying it in scenarios such as e-commerce. Typically, the population is split into two random groups, and the treatment is assigned to one of them ($T=1$). The difference between their response variable $Y$ (e.g., time spent on the system) quantifies the average treatment effect (ATE) of the randomized control trial (RCT) [44], i.e., $Y_{T}-Y_{C}=\mathbb{E}[Y|T=1]-\mathbb{E}[Y|T=0]$, where $T$ and $C$ refer to the treatment and control groups, respectively. In this setting, however, we risk causing churn on a massive scale if the proposed change is not effective [2]. It is thus more wise to test on a smaller scale and try to extrapolate our findings. In these cases, we can build a model that predicts the outcome without actually intervening, based on the samples’ confounders, e.g., for a user, this could be the purchase history and demographics. The core hypothesis is that given the common assumption of unconfoundedness [10], samples with similar confounders will exhibit similar outcomes under the same treatment. Developing a model that can predict the effect of a treatment on unseen samples is referred to as uplift modeling (UM) [42, 3], mainly stemming from business context applications. The uplift refers to the prediction of the outcome’s change due to the treatment, i.e., the ATE, for a set of samples [12]. The problem is akin to individual treatment effect (ITE) estimation, i.e., predicting the outcome of the same sample with and without treatment (counterfactual) and computing their difference.
@@ 3.1 Uplift Modeling with Graph Neural Networks (UMGNet)
$\text{{loss}}_{t}=\text{CrossEntropy}\big{(}T,\text{Sigmoid}(\mathbf{H}^{\top}_{F}\mathbf{W}^{\top}_{F})\big{)}.$ (8)
@@ 3.2 Active Learning for Uplift GNNs (UMGNet-AL)
Aleatoric uncertainty can be measured using the structure of the graph and the feature distance. Acquisition functions based on diverse criteria, including both types of uncertainties, have proven more effective in node-level tasks [63, 17]. We adopt a similar approach and define $D_{u},\forall u\in U$ to be the degree, $Q_{u}$ is the model’s uncertainty, and $b$ is the budget for new training samples in each iteration. According to the literature [63, 17], we define diversity based on feature clustering. Specifically, we cluster the samples with a $k$-means algorithm in a predefined number of clusters. $C_{u}$ is the set of samples included in the cluster of sample $u$. It is used to balance the number of samples from each cluster in each batch. We calculate the distance $M_{u}\in\mathbb{R}^{n}$ between each sample and its cluster centroid and compute the budget for each cluster based on its relative size and $b$, in $\frac{C_{u}}{n}b\in\mathbb{R}^{n}$. We can cast the problem of choosing a batch to add to the training set as ranking based on $U$, $D$, and $M$, with the first constraint being on the batch size. The second constraint, which represents diversity, pertains to how many samples of a given cluster should be included in the train set in each round. Finally, we want to keep the train set balanced in terms of treatment and control samples, hence the third constraint:
@@
The problem can be solved greedily by ranking and feasibility check, as the cluster constraints to not interact between clusters. We solve it in each iteration of batch selection, while $D$, $C$, and $C_{b}$ are precomputed. In practice, we weigh each of the three criteria in the objective function with a coefficient in $[0,1]$ based on the results of validation.
@@ Appendix 0.A Network as a confounder and as a source for interference
Generally, we can make a distinction between the methods that predict the ATE i) using the network as a confounder [57, 20, 35, 7, 26, 14], ii) assuming interference bias caused by the network [37, 8, 33], iii) separating between confounding and interference bias [36, 9]. That said, all methods are predominantly evaluated based on the predicted ITE’s accuracy, the main difference being that the models addressing interference predict the causal estimate while setting the message passing to zero. Our work falls in the first category and follows the same assumptions as the literature regarding interference [57, 35, 7, 14]. Interference in e-commerce bipartite networks is not studied as extensively as in social networks [55, 28]. Its existence remains unproved by experimental studies, mainly because it relies on the type of recommendation algorithm and the nature of the network.